OpenStudy (anonymous):

+-√(-108)=+-6i√(3) The question is telling to take square root from each side and +-6i√(3) was the answer when I solved it. how do you do this?

OpenStudy (anonymous):

\[\sqrt{-108}=\sqrt{36\times 3\times -1}=\sqrt{36}\sqrt{3}\sqrt{-1}=6\sqrt{3}i\]